OpenStudy (anonymous):

Tell whether each o f the following is true or false. Use your knowledge of properties to help you decide. (7p) 1/7 + -p= 0 . Please help!!

Directrix (directrix):

(7p) 1/7 + -p= 0 True or False ============ (7p) 1/7 + -p = 7p * (1/7) - p = p - p = ? @Hannah_

OpenStudy (anonymous):

false?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

What is two minus two? What is the variable p minus the variable p? If the answer is zero, then true. If the answer is not zero, then false. @Hannah_ What do you think is the answer?

OpenStudy (anonymous):

Its zero haha @Directrix so true!

OpenStudy (anonymous):

thank you(:

Directrix (directrix):

You are welcome. True is correct.